OpenStudy (anonymous):

Place the indicated product in the proper location on the grid. (4x^3 + 7y^3z^4)^2 _ _ _ _ _ _ _ x _ _ _ z _ _ _ _ _ _ z _

OpenStudy (zehanz):

It looks terribly complicated, but in the end it's just the form:\[(a+b)^2=a^2+2ab+b^2 \]So:\[(4x^3 + 7y^3z^4)^2=(4x^3)^2+2*4x^3*7y^3z^4+(7y^3z^4)^2\] I'm afraid I don't quite get what this "grid" of yours has to do with it (shouldn't the middle letter be a "y"?), but if I work it out just a little more, I get:\[16x^6+56x^3y^3z^4+49y^6z^8\] That's all I can make of it...

OpenStudy (anonymous):

16x^6 + 2412x^3 y^7 z^4 + 9y^14 z^8 I figured it out. However, I'll give you best response for helping!

OpenStudy (zehanz):

Glad to be of help!